OpenStudy (anonymous):

Use basic identities to simplify the expression. one divided by cotangent of theta to the second power. + sec θ cos θ

OpenStudy (anonymous):

please help me! I am taking precalc online and don't understand a single thing about basic identities

hartnn (hartnn):

so its \(\huge \dfrac{1}{\cot^2 \theta + \sec \theta \cos \theta} \)

hartnn (hartnn):

firstly, use the identity that sec theta = 1/cos theta so, sec theta cos theta = 1 and do you know what is \(\large 1+\cot^2 \theta = ... ?\)

OpenStudy (anonymous):

only the cot^2theta is the denominator

OpenStudy (anonymous):

no I don't know, I am COMPLETELY lost in this subject

hartnn (hartnn):

okk., so anyways you can use sec theta cos theta = 1 \(\large \dfrac{1}{\cot^2 \theta }+1 \)

hartnn (hartnn):

now 1/ cot theta = tan theta so, \(\large \tan^2 \theta+1 \)

hartnn (hartnn):

and ther's a standard pythagorean identity that \(\large 1+\tan^2 \theta = \sec^2 \theta \) thats it! :)

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

can you walk me through another please?

hartnn (hartnn):

sure

OpenStudy (anonymous):

Simplify the expression. quantity cosecant of x to the power of two times secant of x to the power of two divided by quantity secant of x to the power of two plus cosecant of x to the power of two.

OpenStudy (anonymous):